If the tax on a $110.00 item is $8.50, what is the tax on a $150.00 item if the sales tax charged is proportional?
Found 2 solutions by Mathtut, actuary:
Answer by Mathtut(3670)
Answer by actuary(112)

The set-up for this problem is
8.50/110=x/150
Solve for x by multiplying each side of the equation by 150.
(8.50*150)/110=x
x=11.59 (tax on $150)
I hope that this helps.